OpenStudy (anonymous):

given sinx=-1/4 and 180

OpenStudy (amistre64):

lets find the cos of x..... think we can?

OpenStudy (amistre64):

cos(2x) = cos(x)cos(x) - sin(x)sin(x)

OpenStudy (amistre64):

sin^2(x) + cos^2(x) = 1 cos^2(x) = 1 - (-1/4)^2 cos^2(x) = 1 -1/16 cos^2(x) = 15/16 cos(x) = sqrt(15)/4

OpenStudy (amistre64):

cos(2x) = sqrt(15)sqrt(15) -1(-1) ------------- - -------- 4(4) 4(4)

OpenStudy (amistre64):

cos(2x) = 15 - 1 ------ 16 cos(2x) = 14/16 = 7/8

OpenStudy (anonymous):

oh my gosh, that makes total sense now! i just have trouble picking what identity or formula to use.

OpenStudy (amistre64):

trig is 90% memorizing

OpenStudy (anonymous):

yeahh, tell me about it. but thank you!

OpenStudy (amistre64):

the only time your gonna have to actually worry about this stuff is in calculus classes when you try to manipulate formula into derivatives and integrals.... other than that....pointless. A calculator gives you all the answers

OpenStudy (anonymous):

ohh gotcha. but why do you use cos2x= cosxcosx - sinxsinx? i thought that was a sum and difference formula? im just a little confused with that part.

OpenStudy (amistre64):

since they want you to find cos(2x) and give you sin(x) its only natural :)

OpenStudy (anonymous):

oh, okay got it!